OpenStudy (anonymous):

can someone help me? Mortgage is $740 monthly was 59 days late bank charged 15.00 what was the rate is interest assuming simple interest. round to nearest hundredth

OpenStudy (anonymous):

SI = (P*R*T)/100 Therefore R = (SI * 100)/(P*T) P = $740 SI = 15 T = 59/30 (Assuming 30-day months) This will solve for the monthly interest rate.

OpenStudy (anonymous):

not enough info. need the loan amount, the term. with the information given it is around 2%.